A company pays 15,40 euro/ hour during the daytime and 19,60 euro/hour during night time.A man has got 852,60 euro after his work of a certain time during daytime and a certain time during night time. If he changes his night time with that of the day time he will be recieved 25,20 euro less. Calculate how many hors he worked during daytime and night time ?
Answer by checkley71(8403)

15.40X+19.60Y=852.60
15.40X=852.60-19.60Y
X=(852.60-19.60Y)/15.40 AND
15.40Y+19.60X=852.60-25.20
IF WE SUBSTITUTE (852.60-19.60Y)/15.40 FOR X IN THE SECOND EQUATION WE GET
15.40Y+19.60(852.60-19.60Y)/15.40=852.60-52.20
15.40Y+19.60/15.40(852.60-19.60Y)=800.4
15.40Y+1.2727(852.60-19.60Y)=800.4
15.40Y+1085.1272-24.94545Y=800.4
-9.54545Y=800.4-1085.1272
-9.54545Y=-284.7272
Y=-284.7272/-9.54545
Y=29.829 THUS THE NIGHT TIME HOURS WORKED
X=(852.6-19,60*29.829)/15.40
X=(852.6-584.648)/15.40
X=267.952/15.40
X=17.4 THUS THE DAY TIME HOURS WORKED